how long would it take for your money to triple in value if it earned a return of 10%?
Answer #1

Future value=(Present value)*(1+interest rate)^n
Here, n is the time period
Given that the future value=3*Present value and the interest rate is 10%
Substituting the values, we get;
3*(Present value)=(Present value)*(1+10%)^n
3=(1+10%)^n
3=(1.1)^n
Taking natural log on both the sides, we get
Ln(3)=n*Ln(1.1)
=>Ln(3)/Ln(1.1)=n
=>1.098612289/0.09531018=n
=>n=11.52670459 or 11.53 years (Rounded to two decimal places)
So, the time period is 11.53 years